# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ A.1.3 Additional Annotator Evaluation Results
Table 7: Detailed comparison of our annotator against Face++ [42] and InsightFace [43] on VGGFace2 dataset. ‘All’ denotes the averaged metrics across three levels of sample difficulty: ‘Easy,’ ‘Medium,’ and ‘Hard.’ Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings and testings within each level are reported. The best results are shown in Bold.
Level Method A-FF++ [80]
Female Male Young Middle_Aged Senior
precision recall F1 precision recall F1 precision recall F1 precision recall F1 precision recall F1
90.106 88.345 89.129 88.560 90.007 89.201 74.707 72.839 73.547 52.934 81.935 63.981 84.769 37.933 51.193
Face++ [42] (0.282) (0.400) (0.220) (0.323) (0.295) (0.188) (1.264) (0.581) (0.656) (0.790) (1.288) (0.852) (0.922) (1.705) (1.884)
87.918 81.284 84.344 82.787 88.662 85.533 67.148 83.284 74.317 48.659 74.220 58.725 93.455 20.959 33.854
InsightFace [43] (0.458) (0.861) (0.560) (0.625) (0.472) (0.399) (1.429) (0.427) (0.959) (0.658) (1.532) (0.831) (1.228) (1.708) (2.315)
89.851 94.149 91.866 93.177 88.451 90.641 90.888 87.477 88.689 88.461 71.439 78.337 93.390 66.519 77.023
All Ours (0.290) (0.302) (0.202) (0.280) (0.384) (0.226) (0.632) (0.280) (0.396) (1.556) (1.064) (1.012) (0.686) (1.461) (1.138)
98.157 97.947 98.052 97.950 98.159 98.054 85.686 93.398 89.372 69.170 89.520 78.036 95.646 58.880 72.888
Face++ [42] (0.302) (0.136) (0.146) (0.131) (0.308) (0.151) (0.625) (1.072) (0.676) (0.878) (0.431) (0.651) (0.371) (0.546) (0.427)
97.004 96.640 96.820 96.656 97.013 96.833 76.172 98.120 85.762 61.246 86.840 71.828 98.106 28.800 44.524
InsightFace [43] (0.470) (0.605) (0.387) (0.579) (0.482) (0.380) (1.303) (0.665) (1.016) (0.576) (1.039) (0.519) (0.755) (0.633) (0.718)
97.987 99.920 98.944 99.919 97.947 98.923 91.450 100.000 95.530 99.494 94.320 96.838 95.154 85.022 89.802
Easy Ours (0.222) (0.065) (0.092) (0.067) (0.233) (0.097) (0.862) (0.000) (0.470) (0.170) (0.483) (0.286) (0.500) (0.724) (0.478)
98.839 89.590 93.985 90.605 98.960 94.597 84.212 80.092 82.098 48.990 86.710 62.604 91.594 25.400 39.746
Face++ [42] (0.312) (0.691) (0.294) (0.553) (0.285) (0.230) (0.742) (0.379) (0.333) (0.591) (0.585) (0.520) (0.670) (1.688) (2.102)
95.655 79.813 87.017 82.685 96.373 89.005 69.090 85.800 76.542 45.982 73.680 56.622 96.654 15.040 26.022
InsightFace [43] (0.499) (0.798) (0.546) (0.573) (0.433) (0.408) (0.862) (0.358) (0.579) (0.273) (0.985) (0.435) (0.636) (0.794) (1.194)
98.907 98.827 98.866 98.828 98.907 98.867 96.968 97.748 97.356 95.636 79.488 86.812 96.102 63.732 76.632
Medium Ours (0.227) (0.131) (0.103) (0.128) (0.229) (0.104) (0.311) (0.230) (0.158) (0.900) (0.673) (0.415) (0.399) (1.483) (1.134)
73.323 77.498 75.352 77.123 72.901 74.952 54.224 45.026 49.172 40.642 69.576 51.302 67.066 29.518 40.946
Face++ [42] (0.231) (0.374) (0.220) (0.284) (0.292) (0.184) (2.425) (0.291) (0.960) (0.901) (2.847) (1.385) (1.727) (2.881) (3.122)
71.096 67.400 69.194 69.020 72.600 70.762 56.182 65.932 60.648 38.748 62.140 47.724 85.604 19.036 31.016
72.658 83.700 77.787 80.784 68.500 74.133 84.246 64.684 73.180 70.252 40.508 51.360 88.914 50.804 64.634
Hard Ours (0.419) (0.710) (0.411) (0.645) (0.691) (0.477) (0.722) (0.610) (0.561) (3.600) (2.035) (2.334) (1.160) (2.175) (1.803)
Table 8: Detailed comparison of our annotator against Face++ [42] and InsightFace [43] on A-FF++ dataset. ‘All’ denotes the averaged metrics across three levels of sample difficulty: ‘Easy,’ ‘Medium,’ and ‘Hard.’ Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings and testings within each level are reported. The best results are shown in Bold.
Level Method A-DFDC [19]
Female Male Young Middle_Aged Senior
precision recall F1 precision recall F1 precision recall F1 precision recall F1 precision recall F1
74.751 83.324 78.453 78.824 68.797 72.922 74.499 81.417 77.479 58.696 74.613 65.608 85.533 52.741 63.542
Face++ [42] (0.675) (0.948) (0.603) (0.933) (0.929) (0.732) (0.567) (1.186) (0.772) (0.688) (0.420) (0.496) (0.792) (1.559) (1.469)
72.691 64.662 68.376 68.160 75.560 71.623 71.209 86.658 77.985 58.102 76.777 66.086 83.061 36.659 49.620
InsightFace [43] (0.649) (1.208) (0.671) (0.779) (0.754) (0.453) (0.852) (0.703) (0.676) (1.040) (0.370) (0.797) (0.989) (1.802) (1.833)
74.805 88.720 80.839 85.137 67.864 74.834 88.669 76.413 81.695 91.753 77.845 83.772 90.797 77.147 82.823
All Ours (0.588) (0.587) (0.498) (0.622) (0.848) (0.619) (0.633) (1.103) (0.752) (0.736) (0.670) (0.545) (0.618) (1.356) (1.116)
94.124 89.917 91.965 90.351 94.367 92.309 86.586 96.596 91.316 73.340 85.560 78.974 99.440 71.324 83.066
Face++ [42] (0.977) (1.010) (0.564) (0.835) (1.025) (0.535) (0.691) (0.713) (0.489) (0.853) (0.933) (0.622) (0.255) (1.013) (0.768)
91.874 78.693 84.753 81.389 93.013 86.801 73.580 95.760 83.216 62.690 77.960 69.496 94.034 42.760 58.776
InsightFace [43] (0.986) (1.798) (0.744) (1.140) (1.059) (0.384) (0.821) (0.196) (0.552) (0.659) (0.898) (0.746) (1.262) (1.209) (1.189)
95.289 97.413 96.337 97.353 95.173 96.248 97.448 99.240 98.334 94.902 96.996 95.938 99.730 89.844 94.528
Easy Ours (1.019) (0.136) (0.566) (0.157) (1.098) (0.621) (0.635) (0.233) (0.359) (0.030) (0.540) (0.280) (0.168) (0.433) (0.303)
70.438 81.415 75.529 77.764 65.533 71.124 70.980 64.526 67.584 53.830 69.930 60.828 73.342 58.332 64.972
Face++ [42] (0.334) (0.881) (0.528) (0.836) (0.418) (0.457) (0.703) (2.147) (1.379) (0.578) (0.000) (0.372) (0.815) (1.562) (1.158)
69.126 66.733 67.905 67.859 70.200 69.006 73.794 77.866 75.768 54.900 71.330 62.038 68.248 44.000 53.474
InsightFace [43] (0.316) (1.285) (0.813) (0.788) (0.221) (0.345) (1.105) (1.205) (0.966) (1.229) (0.000) (0.780) (0.711) (2.241) (1.769)
69.072 95.067 80.011 92.097 57.433 70.746 89.568 67.934 77.250 94.220 69.246 79.824 82.306 82.014 82.158
Medium Ours (0.350) (0.680) (0.444) (1.027) (0.512) (0.593) (0.325) (2.048) (1.451) (1.257) (0.431) (0.524) (1.062) (0.908) (0.891)
59.692 78.639 67.866 68.357 46.489 55.334 65.930 83.128 73.536 48.918 68.350 57.022 83.816 28.568 42.588
Face++ [42] (0.715) (0.953) (0.716) (1.128) (1.344) (1.203) (0.307) (0.698) (0.449) (0.635) (0.326) (0.495) (1.307) (2.102) (2.481)
57.073 48.560 52.471 55.231 63.467 59.062 66.252 86.348 74.972 56.716 81.042 66.724 86.900 23.216 36.610
InsightFace [43] (0.647) (0.543) (0.456) (0.409) (0.983) (0.629) (0.631) (0.707) (0.511) (1.233) (0.211) (0.864) (0.995) (1.955) (2.540)
60.054 73.680 66.170 65.961 50.987 57.509 78.990 62.066 69.500 86.136 67.294 75.554 90.354 59.582 71.784
Hard Ours (0.394) (0.945) (0.485) (0.683) (0.934) (0.645) (0.940) (1.029) (0.447) (0.922) (1.038) (0.832) (0.624) (2.726) (2.154)
Table 9: Detailed comparison of our annotator against Face++ [42] and InsightFace [43] on A-DFDC dataset. ‘All’ denotes the averaged metrics across three levels of sample difficulty: ‘Easy,’ ‘Medium,’ and ‘Hard.’ Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings and testings within each level are reported. The best results are shown in Bold.
Level Method A-DFD[19]
@@
precision recall F1 precision recall F1
74.375 62.743 67.925 68.258 78.854 73.096
Face++ [42] (1.442) (1.445) (1.256) (1.197) (1.544) (1.228)
71.967 51.796 59.975 63.600 81.636 71.405
InsightFace [43] (1.062) (1.161) (1.053) (0.651) (0.863) (0.660)
72.884 83.547 77.548 78.938 66.418 71.615
All Ours (0.580) (0.753) (0.538) (0.750) (0.979) (0.783)
95.014 82.284 88.179 84.398 95.677 89.676
Face++ [42] (0.377) (1.876) (1.011) (1.378) (0.391) (0.686)
94.405 75.573 83.944 79.639 95.520 86.858
InsightFace [43] (0.777) (0.956) (0.794) (0.683) (0.634) (0.596)
94.434 93.600 94.013 93.659 94.480 94.066
Easy Ours (0.418) (0.566) (0.307) (0.517) (0.451) (0.295)
65.536 60.151 62.715 63.114 68.283 65.587
Face++ [42] (1.779) (1.249) (1.202) (1.077) (2.356) (1.566)
64.397 47.147 54.434 58.323 73.947 65.211
InsightFace [43] (1.071) (1.433) (1.310) (0.769) (0.646) (0.671)
65.158 86.133 74.188 79.536 53.920 64.258
Medium Ours (0.886) (0.625) (0.666) (0.911) (1.776) (1.472)
62.576 45.793 52.882 57.261 72.603 64.024
Face++ [42] (2.170) (1.210) (1.556) (1.137) (1.885) (1.433)
57.100 32.667 41.547 52.838 75.440 62.145
InsightFace [43] (1.339) (1.096) (1.054) (0.501) (1.310) (0.714)
59.062 70.907 64.442 63.618 50.853 56.520
Hard Ours (0.437) (1.068) (0.641) (0.820) (0.709) (0.582)
Table 10: Detailed comparison of our annotator against Face++ [42] and InsightFace [43] on A-DFD dataset. ‘All’ denotes the averaged metrics across three levels of sample difficulty: ‘Easy,’ ‘Medium,’ and ‘Hard.’ Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings and testings within each level are reported. The best results are shown in Bold.
A-Celeb-DF-v2 [19]
Female Male
Level Method precision recall F1 precision recall F1
97.624 83.302 89.455 86.517 98.318 91.819
Face++ [42] (0.815) (0.727) (0.438) (0.494) (0.640) (0.453)
97.442 85.842 91.041 88.079 98.007 92.635
InsightFace [43] (0.537) (0.569) (0.317) (0.388) (0.464) (0.299)
96.381 93.611 94.921 94.147 96.687 95.355
All Ours (0.756) (0.564) (0.400) (0.422) (0.782) (0.426)
99.889 96.480 98.155 96.598 99.893 98.218
Face++ [42] (0.055) (0.418) (0.236) (0.392) (0.053) (0.222)
99.837 98.107 98.964 98.140 99.840 98.983
InsightFace [43] (0.054) (0.352) (0.180) (0.340) (0.053) (0.174)
100.000 99.973 99.987 99.973 100.000 99.987
Easy Ours (0.000) (0.053) (0.027) (0.053) (0.000) (0.027)
99.732 89.227 94.185 90.260 99.760 94.771
Face++ [42] (0.199) (0.952) (0.558) (0.787) (0.177) (0.460)
99.639 88.320 93.638 89.513 99.680 94.323
InsightFace [43] (0.226) (0.496) (0.354) (0.412) (0.200) (0.298)
99.760 99.760 99.760 99.760 99.760 99.760
Medium Ours (0.053) (0.177) (0.100) (0.176) (0.053) (0.100)
93.251 64.200 76.024 72.694 95.300 82.469
Face++ [42] (2.190) (0.812) (0.519) (0.304) (1.691) (0.679)
92.850 71.100 80.521 76.584 94.500 84.600
InsightFace [43] (1.331) (0.860) (0.417) (0.412) (1.140) (0.424)
89.384 81.100 85.015 82.707 90.300 86.319
Hard Ours (2.214) (1.463) (1.073) (1.037) (2.294) (1.150)
@@
Table 11: Detailed comparison of our annotator against Face++ [42] and InsightFace [43] on A-Celeb-DF-v2 dataset. ‘All’ denotes the averaged metrics across three levels of sample difficulty: ‘Easy,’ ‘Medium,’ and ‘Hard.’ Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings and testings within each level are reported. The best results are shown in Bold.
Level Method Adience [81]
Female Male Young Middle_Aged Senior
precision recall F1 precision recall F1 precision recall F1 precision recall F1 precision recall F1
71.800 87.124 76.457 84.701 71.257 76.322 89.823 68.558 77.632 51.442 70.678 58.668 50.280 76.456 60.540
Face++ [42] (0.900) (0.743) (0.797) (0.732) (0.863) (0.601) (0.444) (1.024) (0.696) (1.569) (1.397) (1.313) (3.051) (3.228) (2.199)
67.003 74.180 68.487 73.848 67.877 69.512 86.510 37.559 51.425 37.028 65.519 45.545 26.775 72.412 38.946
InsightFace [43] (0.687) (1.140) (0.650) (1.068) (0.722) (0.583) (1.147) (1.079) (1.225) (0.888) (1.304) (0.884) (1.672) (2.321) (2.045)
78.103 94.416 83.852 96.915 82.496 88.602 77.517 81.767 79.578 48.206 37.327 41.843 67.352 69.370 68.241
All Ours (0.732) (0.338) (0.603) (0.382) (0.607) (0.443) (0.500) (0.963) (0.579) (1.794) (0.529) (1.061) (3.613) (3.679) (3.258)
97.977 93.613 95.658 92.325 97.319 94.754 96.754 81.939 88.730 33.469 68.722 44.960 59.861 88.473 71.165
Face++ [42] (0.251) (0.646) (0.273) (0.660) (0.344) (0.284) (0.253) (0.746) (0.434) (2.274) (1.985) (2.095) (4.706) (4.331) (2.485)
96.613 89.815 93.087 88.180 96.009 91.925 97.076 58.063 72.662 19.981 70.808 31.150 26.572 84.972 40.411
InsightFace [43] (0.210) (1.000) (0.537) (0.980) (0.331) (0.540) (0.455) (0.819) (0.717) (0.942) (2.276) (1.185) (2.943) (2.722) (3.521)
99.822 99.974 99.898 99.969 99.776 99.872 93.651 97.159 95.372 57.920 37.414 45.375 97.721 96.831 97.221
Easy Ours (0.104) (0.052) (0.078) (0.063) (0.124) (0.094) (0.599) (0.507) (0.455) (2.428) (0.217) (1.469) (3.063) (2.986) (2.037)
82.280 87.626 84.863 72.271 63.091 67.350 90.557 64.622 75.413 57.229 71.100 63.397 47.362 75.972 58.295
Face++ [42] (1.064) (0.710) (0.617) (1.249) (1.654) (0.996) (0.379) (1.444) (1.015) (1.735) (1.580) (1.350) (2.070) (3.031) (1.800)
78.900 75.366 77.087 55.589 60.493 57.923 83.417 30.372 44.510 40.618 61.066 48.776 26.303 65.071 37.453
InsightFace [43] (1.250) (0.827) (0.818) (0.852) (1.375) (0.637) (2.029) (2.139) (2.587) (0.934) (0.713) (0.618) (1.300) (2.274) (1.616)
92.671 99.184 95.816 98.159 84.639 90.897 76.159 80.899 78.455 40.740 34.599 37.409 69.317 69.799 69.532
Medium Ours (0.607) (0.280) (0.385) (0.599) (0.735) (0.505) (0.372) (0.913) (0.492) (1.307) (0.808) (0.812) (5.848) (5.737) (5.645)
35.144 80.134 48.851 89.507 53.361 66.860 82.159 59.114 68.752 63.628 72.212 67.646 43.617 64.923 52.162
Face++ [42] (1.384) (0.874) (1.502) (0.287) (0.591) (0.522) (0.701) (0.882) (0.638) (0.697) (0.626) (0.495) (2.376) (2.322) (2.311)
25.495 57.358 35.287 77.776 47.129 58.688 79.037 24.242 37.102 50.485 64.681 56.708 27.452 67.194 38.973
InsightFace [43] (0.602) (1.594) (0.596) (1.373) (0.460) (0.572) (0.956) (0.280) (0.372) (0.789) (0.923) (0.848) (0.773) (1.967) (0.999)
41.818 84.089 55.842 92.619 63.072 75.038 62.742 67.243 64.907 45.958 39.968 42.745 35.018 41.479 37.972
Hard Ours (1.484) (0.684) (1.345) (0.486) (0.962) (0.730) (0.530) (1.468) (0.790) (1.648) (0.563) (0.902) (1.929) (2.314) (2.092)
Table 12: Detailed comparison of our annotator against Face++ [42] and InsightFace [43] on Adience dataset. ‘All’ denotes the averaged metrics across three levels of sample difficulty: ‘Easy,’ ‘Medium,’ and ‘Hard.’ Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings and testings within each level are reported. The best results are shown in Bold.
A.2 Phase2: Demographically Annotation Generation
A.2.1 Detailed Information of Datasets
Methods #Samples FFHQ CASIA-WebFace IMDB-WIKI CelebA A-FF+ A-DFDC A-DFD A-Celeb-DF-v2
 [6]  [37]  [38]  [36] (Real) [80] (Real)  [80] (Real) [80] (Real) [80]
A-FF++  [2] 105K
A-DFDC  [39] 37K
A-DFD  [40] 31K
A-Celeb-DF-v2  [41] 155K
AttGAN [49] 6K
MMDGAN [50] 1K
StarGAN [49] 5.6K
StyleGAN [49] 10K
StyleGAN2 [51] 118K
StyleGAN3 [52] 26.7K
MSG-StyleGAN [50] 1K
ProGAN [53] 100K
STGAN [50] 1K
VQGAN [54] 50K
DALLE2 [55] 204
IF [55] 505
Midjourney [55] 100
DCFace [56] 529K
Latent Diffusion [57] 20K
Palette [58] 6K
SD v1.5 [59] 18K
SD Inpainting [59] 20.9K
Total 1,245,660 70,000 474,876 26,788 202,502 21,593 37,836 8,856 23,645
866,096
Table 13: Number of real and fake images from different fake image datasets and their corresponding real image sources.
Table 13 shows the detailed information of all subsets we collected and incorporated into our AI-Face dataset. It covers fake facial images from deepfake videos, generated from GANs and DMs. The corresponding real sources of most AI-generated face subsets are FFHQ [6] and CelebA [36]. In general, our AI-Face dataset contains 30 subsets (22 fake subsets and 8 real subsets) and 37 generation methods ( methods are summed as 5 in A-FF++, 5 in A-DFD, 8 in A-DFDC, 1 in A-Celeb-DF-v2, 10 GANs, and 8 DMs), including a total of 1,245,660 fake face images and 866,096 real face images. Fig. A.1 visualizes face images of each subset. Fig. A.2 further demonstrates the detailed demographic distribution of our AI-Face dataset. The dataset is relatively gender-balanced, and the subjects are majorly young and white individuals.
Figure A.1: Visualization of images in AI-Face dataset. SD is short for Stable Diffusion.
Figure A.2: Detailed demographic distribution of AI-Face dataset.
A.2.2 Details of Threshold Settings for Human Correction
In this section, we present uncertainty score distributions of each attribute (i.e., Gender, Age, and Race) of each subset in our AI-Face dataset, as shown from Fig. A.4 to Fig. A.31. Overall, our annotator shows higher confidence in predicting gender attributes compared to predicting age, as observed from these uncertainty score distributions. It is clear that different subsets show different distributions, so we dynamically adjust the threshold $t^{{a}_{j}}$ for each attribute $a$ on subset $j$ defined in ‘Human Correction’ in Section 3.2. First, we fit the distribution with gamma distribution and calculate its mean and standard deviation. Then, the $t^{{a}_{j}}$ is calculated using $Mean+\lambda Std$. After getting the threshold, we can get the total image number within each subset that needed human correction. We assume it takes three seconds for a human to correct one annotation for one image, then we can calculate the total time needed for a human to correct these images beyond the threshold. Therefore, The $\lambda$ is dynamically adjusted based on the distribution and the total time needed for human correction.
A.2.3 Examples of Mislabeled images in A-FF++ and A-DFDC
In the evaluation results for Q1 in Section 3.2, we have validated that we cannot directly incorporate existing annotations into our AI-Face dataset. Fig. A.3 displays some image examples where annotations in A-FF++ [80] and A-DFDC [80] are inconsistent with the annotations given by our annotator. A-FF and A-DFDC have mislabeled annotations for ambiguous facial images, whereas our annotator can accurately predict them. This visualization of images further validates that existing annotations cannot be directly merged into our dataset.
A.2.4 Additional Results of Validating the Effectiveness of Human Correction Strategy
Since A-DFD [19] and A-Celeb-DF-v2 [19] provide gender annotation, we can compare our two versions of datasets with it. One is Ours before human correction (i.e., Ours-DFD(w/o Correction) and Ours-A-Celeb-DF-v2(w/o Correction)), another one is Ours after human correction (i.e., Ours-DFD (Correction) and Ours-A-Celeb-DF-v2 (Correction)). As same setting as in evaluation for Q2 Section 3.2, we sample 1,200 attribute-balanced images (400 easy, 400 medium, and 400 hard) based on uncertainty score. Three humans re-annotated these images to establish ground truth. As shown in Table 14, Ours-DFD (Correction) and Ours-A-Celeb-DF-v2 (Correction) outperforms ours without correction version and A-DFD and A-Celeb-DF-v2 (e.g., the accuracy of Ours-DFD (Correction) is 22.866% higher than A-DFD and 13.526% higher than Ours-DFD(w/o Correction)). This suggests that our dataset annotation quality is much better than the existing annotation in A-DFD [19] and A-Celeb-DF-v2 [19]. And our human correction strategy further improves our dataset annotation quality.
Gender
ACC Precision Recall F1
A-DFD [19] 70.612 71.347 74.245 69.900
Ours-DFD(w/o Correction) 79.952 79.868 83.979 79.308
Ours-DFD (Correction) 93.478 91.673 95.034 92.898
A-Celeb-DF-v2 [19] 89.697 90.622 90.622 89.697
Ours-A-Celeb-DF-v2(w/o Correction) 91.414 91.404 91.831 91.391
Ours-A-Celeb-DF-v2 (Correction) 93.535 93.655 94.087 93.525
Table 14: Evaluation results to demonstrate the effectiveness of human correction strategy.
Figure A.3: Examples of face images where annotations in A-FF++ and A-DFDC are inconsistent with Ours-FF++ (w/o Correction) and Ours-DFDC (w/o Correction).
Figure A.4: Uncertainty Score histograms for gender, age and race for A-FF++  [2] dataset.
Figure A.5: Uncertainty Score histograms for gender, age and race for A-DFDC [39] dataset.
Figure A.6: Uncertainty Score histograms for gender, age and race for A-DFD [40] dataset.
Figure A.7: Uncertainty Score histograms for gender, age and race for A-Celeb-DF-v2 [41] dataset.
Figure A.8: Uncertainty Score histograms for gender, age and race for AttGAN [49] dataset.
Figure A.9: Uncertainty Score histograms for gender, age and race for MMDGAN [50] dataset.
Figure A.10: Uncertainty Score histograms for gender, age and race for StarGAN [49] dataset.
Figure A.11: Uncertainty Score histograms for gender, age and race for StyleGAN [49] dataset.
Figure A.12: Uncertainty Score histograms for gender, age and race for StyleGAN2  [51] dataset.
Figure A.13: Uncertainty Score histograms for gender, age and race for StyleGAN3 [52] dataset.
Figure A.14: Uncertainty Score histograms for gender, age and race for MSG-StyleGAN [50] dataset.
Figure A.15: Uncertainty Score histograms for gender, age and race for ProGAN [53] dataset.
Figure A.16: Uncertainty Score histograms for gender, age and race for STGAN [50] dataset.
Figure A.17: Uncertainty Score histograms for gender, age and race for VQGAN [54] dataset.
Figure A.18: Uncertainty Score histograms for gender, age and race for DALLE2 [55], IF [55], Midjourney [55] dataset.
Figure A.19: Uncertainty Score histograms for gender, age and race for DCFace [56] dataset.
Figure A.20: Uncertainty Score histograms for gender, age and race for Latent Diffusion [57] dataset.
Figure A.21: Uncertainty Score histograms for gender, age and race for Palette [58] dataset.
Figure A.22: Uncertainty Score histograms for gender, age and race for SD v1.5 [59] dataset.
Figure A.23: Uncertainty Score histograms for gender, age and race for SD Inpainting  [59] dataset.
Figure A.24: Uncertainty Score histograms for gender, age and race for FFHQ [6] dataset.
Figure A.25: Uncertainty Score histograms for gender, age and race for CASIA-WebFace [37] dataset.
Figure A.26: Uncertainty Score histograms for gender, age and race for IMDB-WIKI [38] dataset.
Figure A.27: Uncertainty Score histograms for gender, age and race for CelebA [36] dataset.
Figure A.28: Uncertainty Score histograms for gender, age and race for A-FF++ [2] real dataset.
Figure A.29: Uncertainty Score histograms for gender, age and race for A-DFDC [39] real dataset.
Figure A.30: Uncertainty Score histograms for gender, age and race for A-DFD [40] real dataset.
Figure A.31: Uncertainty Score histograms for gender, age and race for A-Celeb-DF-v2 [41] real dataset.
Appendix B Fairness Benchmark
B.1 Details of Detection Methods

### v3

@@ B.3 Experimental Study of Existing Face Attribute Prediction Tools
Table B.2: Detailed comparison of our annotator against Face++ [71] and InsightFace [72] on IMDB-WIKI  [20] dataset. Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings are reported. The best results are shown in Bold.
@@
Dataset Method precision recall F1 precision recall F1
0.9816 0.9795 0.9805 0.9795 0.9816 0.9805
Face ++ (0.3021) (0.1360) (0.1459) (0.1312) (0.3084) (0.1508)
0.9700 0.9664 0.9682 0.9666 0.9713 0.9683
Insightface (0.4697) (0.6046) (0.3867) (0.5794) (0.4815) (0.3802)
0.9799 0.9992 0.9894 0.9992 0.9795 0.9892
FF++ Ours (0.0022) (0.0006) (0.0009) (0.0007) (0.0023) (0.0009)
0.9412 0.8992 0.9197 0.9035 0.9437 0.9231
Face ++ (0.9771) (1.0095) (0.5639) (0.8353) (1.0246) (0.5353)
0.9187 0.7869 0.8475 0.8139 0.9301 0.8680
Insightface (0.9855) (1.7976) (0.7444) (1.1401) (1.0587) (0.3842)
0.9519 0.9741 0.9629 0.9735 0.9507 0.0114
DFDC Ours (0.0106) (0.0014) (0.0059) (0.0015) (0.9619) (0.0064)
0.9501 0.8228 0.8818 0.8440 0.9568 0.8968
Face ++ (0.3773) (1.8758) (1.0113) (1.3784) (0.3907) (0.6856)
0.9441 0.7557 0.8394 0.7964 0.9552 0.8686
Insightface (0.7765) (0.9555) (0.7937) (0.6834) (0.6344) (0.5961)
0.9378 0.9365 0.9366 0.9366 0.9379 0.9372
DFD Ours (0.0045) (0.0053) (0.0033) (0.0049) (0.0048) (0.0033)
0.9989 0.9648 0.9815 0.9660 0.9989 0.9822
Face ++ (0.0553) (0.4182) (0.2361) (0.3918) (0.0533) (0.2215)
0.9984 0.9811 0.9896 0.9814 0.9984 0.9898
Insightface (0.0541) (0.3518) (0.1801) (0.3396) (0.0534) (0.1737)
1.0000 0.9997 0.9999 0.9997 1.0000 0.9999
Celeb-DF-v2 Ours (0.0000) (0.0005) (0.0003) (0.0005) (0.0000) (0.0003)
@@
Table B.3: Detailed comparison of our annotator against Face++ [71] and InsightFace [72] on FF++ [2], DFDC [21], DFD [22], and Celeb-DF-v2 [23] datasets. Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings. The best results are shown in Bold.
B.4 Annotator Implementation Detail
[…]
Figure B.3: Visualization of the skin tone estimation process.
B.5 Details of Human Labeling Activities in Annotation Quality Assessment
The annotation process for assessing the quality of AI-generated face image annotations followed a structured and ethically grounded methodology. Prior to labeling, all human annotators signed an Annotator Agreement outlining the project objectives, confidentiality requirements, and detailed labeling guidelines for gender and age classification. This agreement emphasized impartiality, respect, and adherence to professional conduct throughout the annotation activities. Human annotators then underwent tutorial training using real example images to familiarize themselves with demographic attributes and labeling criteria, focusing on identifying gender-specific features, such as facial structure and presence of facial hair, and age-related indicators, including wrinkles and skin elasticity.
Following the agreement and training, human annotators independently labeled the images based on the established criteria, categorizing gender and age into predefined classes, and recorded their annotations in a CSV file. A structured conflict resolution approach ensured accuracy and consistency in annotations. Labels agreed upon by a majority of annotators were finalized directly, while unanimous disagreements were resolved through collaborative discussions guided by the annotation guidelines. This process ensured that all annotations were objective, reliable, and aligned with ethical standards set forth in the signed agreement.
B.6 Visualization of Skin Tone Annotation Generation
The visualization shown in Fig. B.3 illustrates the skin tone estimation process using the Monk Skin Tone (MST) Scale. Each row represents a sample image, showing the progression from the original face with facial landmarks to the masked skin region that excludes non-skin areas like eyes and lips. Subsequently, the K-Means clustered skin region highlights the dominant skin tones extracted from the facial area. On the right, bar plots display the proportions of the top three dominant tones within the clustered region, with the top tone (largest cluster) mapped to the closest MST Scale shade. This mapping is achieved by calculating the maximum similarity, as indicated by the Euclidean distance in RGB space between the cluster centroid and MST reference colors. This process visually demonstrates how the methodology isolates, clusters, and estimates skin tones for accurate skin tone annotation generation.
Appendix C Fairness Benchmark Settings
C.1 Implementation Detail
For fairness benchmark, all experiments are based on the PyTorch with a single NVIDIA RTX A6000 GPU. During training, we utilize SGD optimizer with a learning rate of 0.0005, with momentum of 0.9 and weight decay of 0.005. The batch size is set to 128 for most detectors. However, for the SRM [87], UCF [26], and PG-FDD [30], the batch size is adjusted to 32 due to GPU memory. For hyperparameters defined in these detectors, we use the default values set in their original papers. All detectors are initialized with their official pre-trained weights, and trained for 10 epochs.
C.2 Details of Detection Methods
We summarized the backbone architecture, GitHub repository link, and publication venue of the detectors implemented in our fairness benchmark in Table C.1. A brief introduction to each detector is provided below: